Fermi National Laboratory
 

Operation Summary for November 2, 2011 - November 4, 2011

 

Wednesday November 2

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI) , and with the MiniBooNE and NuMI experiments taking beam .

During the day shift, operators reported that the New Muon Lab (NML) started a scheduled 15-minute power outage at 9:24 AM. Operators reported that MTA started a 40-minute scheduled power outage at 9:39 AM. A Linac person began a 30-minute pulse study at 10:20 AM. A Booster person began a pulse study at 10:38 AM. A MI person began a g minus 2 (g-2) study at 1:54 PM. MI personnel began a Recycler BPM study at 1:57 PM.

During the evening shift, the g-2 study ended at 4:32 PM. The Recycler BPM study ended.

Thursday November 3

The midnight shift began with Operations monitoring beam to the NuMI and MiniBooNE experiments .

During the midnight shift, at 7:16 AM, a Booster safety system detector alarmed high and tripped off; operators reset the system. A minute later the Booster safety system alarmed again, but this time due to a missed pulse; a Safety System person changed out the detector.

During the day shift, at 8:30 AM, Neutron Therapy Facility personnel began preparing to treat one patient later in the morning. A Booster person began a partial batch kicker study at 3:52 PM. MI personnel began a Recycler BPM study. Operators reported short downtimes due to a Linac RF station (LRF5).

During the evening shift, the Booster kicker study ended at 4:48 PM. The Recycler BPM study ended. At 7:39 PM, two I- Source vacuum pump power supplies tripped off; they reset. The NuMI beam permit tripped off at 8:34 PM due to a MI RF station (MIRF16) trip; everything reset. At 9:55 PM, the Preaccelerator temperature alarmed when the temperature dropped 20 degrees in an instant; about an hour later the temperature jumped back to normal. Operators searched and secured MTA at 11:38 PM.

Friday November 4

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE experiments.

During the midnight shift, operators established beam to the Switchyard (SY) dump at 6:49 AM. A SY person began a study at 7:08 AM.

Operations Coordinator (OpsCo) Plans for Today and the Weekend

The plans for today are to keep delivering protons to the experiments.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/04/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory